OpenStudy (anonymous):

graph of -x - 2y = 6.

OpenStudy (firekat97):

firstly you want to convert the function into point gradient form which looks like y = mx + b this will make it a lot easier to plot

OpenStudy (anonymous):

so 2y=-x+6?

OpenStudy (jackthegreatest):

no

OpenStudy (firekat97):

not quite, the -x should become a + x when you take it over to the other side, and you still have to divide both sides by 2 to get the y by itself remember

OpenStudy (jackthegreatest):

x+2y=-6 2y =-x-6

OpenStudy (firekat97):

-2* not 2 sorry

OpenStudy (jackthegreatest):

after 2y=-x-6 y= -1/2x -3

OpenStudy (anonymous):

but then how do i make that into a graph

OpenStudy (jackthegreatest):

one of the coordinates are (0, -3)

OpenStudy (anonymous):

sorry for all the questions

OpenStudy (jackthegreatest):

the slope is -1/2, so u should be able to go on from there

OpenStudy (jackthegreatest):

ill just give u two more points to help u

OpenStudy (jackthegreatest):

-4,2 -5, 4

OpenStudy (jackthegreatest):

wait nvm i dont think those points work

OpenStudy (jackthegreatest):

but do u get the idea? @Acs